OpenStudy (anonymous):

10.1 g CaO is dropped into a styrofoam coffee cup containing 157 g H2O at 18.0°C. If the following reaction occurs, then what temperature will the water reach, assuming that the cup is a perfect insulator and that the cup absorbs only a negligible amount of heat? [specific heat of water = 4.18 J/g·°C] CaO(s) + H2O(l) → Ca(OH)2(s) triangleH^0(subcript rxn) = −64.8 kJ/mol a. 18.02°C b. 35.8°C c. 311°C d. 42.2°C e. 117°C someone explain to me plz

OpenStudy (aaronq):

the heat released by the reaction (system) raises the temperature of the water (surroundings), \(q_{system}=-q_{surroundings}\) you use \(q=m*C_p*\Delta T\); where \(\Delta T= T_f-Ti\) \(q_{water}=m*C_p*(\color{red}{T_f}-Ti)\) you're solving for the final temp, \(T_f\). BUT first, convert 10.1 g CaO to moles, then find the heat released, \(q_{system}\). \(q_{system}=n_{CaO}*\Delta H_{rxn}\); \( \Delta H_{rxn}\)=−64.8 kJ/mol